# HRGR: Enhancing Image Manipulation Detection via Hierarchical Region-aware Graph Reasoning

### v1

@@ Abstract
Image manipulation detection is to identify the authenticity of each pixel in images. One typical approach to uncover manipulation traces is to model image correlations. The previous methods commonly adopt the grids, which are fixed-size squares, as graph nodes to model correlations. However, these grids, being independent of image content, struggle to retain local content coherence, resulting in imprecise detection. To address this issue, we describe a new method named Hierarchical Region-aware Graph Reasoning (HRGR) to enhance image manipulation detection. Unlike existing grid-based methods, we model image correlations based on content-coherence feature regions with irregular shapes, generated by a novel Differentiable Feature Partition strategy. Then we construct a Hierarchical Region-aware Graph based on these regions within and across different feature layers. Subsequently, we describe a structural-agnostic graph reasoning strategy tailored for our graph to enhance the representation of nodes. Our method is fully differentiable and can seamlessly integrate into mainstream networks in an end-to-end manner, without requiring additional supervision. Extensive experiments demonstrate the effectiveness of our method in image manipulation detection, exhibiting its great potential as a plug-and-play component for existing architectures.
1 Introduction
@@
Figure 1: The difference between (a) previous methods and (b) ours. Instead of using regular grids in previous methods, we model image correlations across different scales (orange and purple arrows) using connected local content-coherent feature regions.
To combat this, countermeasures have been proposed and gained increasing attention in recent years. Deep learning-based methods have emerged as the dominant method for detecting manipulations e.g., [2, 37, 13, 17, 21, 34]. Leveraging the impressive learning capabilities of deep learning models, these methods have shown superior performance compared to earlier traditional forensics techniques [22, 15, 23]. Based on deep neural networks, feature enhancement strategies have been explored to further emphasize manipulation traces. Particularly, exploiting image correlations is a recent feature enhancement solution that has proven effective in enhancing detection ability [17, 21]. The intuition is that manipulated regions can be better perceived when compared to pristine regions. To capture the correlations, these methods usually employ the mechanism of either attention mechanism [21, 34] or graphs [26, 17]. For example, PSCC-Net [21] proposed a spatial-channel correlation module to capture the correlations using attention. ObjectFormer [34] learned the correlation between intermediate features in objects and patches through cross-attention. HGCN [26] and ERMPC [17] sought the relationship between pristine and manipulated regions by building a graph structure.
@@
Question: “Could we model the image correlations based on the units that are dependent on the image content rather than the regular grids?”
@@
Definitation: The connected local content-coherent region on feature maps is called Feature Region. These regions are usually irregular and dependent on the image content. Remark: In contrast, the grids are regular and fixed squares that are independent of image content.
@@
2 Related Works
The recent forensics methods mainly rely on deep neural networks to expose the manipulations [37, 13, 11, 3, 21, 34, 16, 41, 30, 43, 8, 42]. ManTra-Net [37], SPAN [13], TransForensics [11], MVSS-Net [3], PSCC-Net [21] and HiFi-Het [8] are representative methods that are designed on specific DNNs. More recently, Vision Transformer (ViT) [5] are also adapted in our task, such as ObjectFormer [34] and EVP [20]. Within these methods, strategies using graph spirits for capturing the image correlations have been studied to further enhance feature effectiveness. HGCN-Net [26] introduced graph structures for manipulation detection tasks, which modeled region adjacency relationships on feature grids by considering four connectivity. PSCC-Net [21] introduced a module for capturing spatial correlations using attention. This module can be conceptualized as a fully connected graph, where each element corresponds to a grid and is connected to every other element. ERMPC [17] extracted the inconsistency between real and manipulated regions by controlling message passing. It is worth highlighting that the aforementioned methods share a commonality in constructing graphs based on grids. However, due to the content independence and fixed sizes of these grids, they lack an understanding of the content coherence and often struggle to capture irregularly manipulated regions. As such, we describe the Hierarchical Region-aware Graph Reasoning method to address these limitations.
3 Method
In contrast to previous works that use grids for correlation modeling, we develop an end-to-end Hierarchical Region-aware Graph Reasoning method to effectively excavate the image correlations. The overview of our method is illustrated in Fig. 2 (left).
Our method consists of two main components: Hierarchical Region-aware Graph and Structure-agnostic Graph Reasoning. The Hierarchical Region-aware Graph is a newly proposed graph structure constructed on content-coherence feature regions and the Structure-agnostic Graph Reasoning passes the messages through the constructed graph to enhance the effectiveness of features. These two components are elaborated in Sec. 3.1 and Sec. 3.2 respectively. Then we describe the training process of our method in Sec. 3.3. The overall process of our method is shown in Alg. 1.
3.1 Hierarchical Region-aware Graph
@@
3.1.1 Differentiable Feature Partition
@@
Intuitively, the conventional clustering algorithms (e.g., K-means) can be applied to feature layers by measuring the similarity among feature elements. However, this algorithm suffers from two difficulties. One is that local connectivity can not be guaranteed and the other is the clustering process is not differentiable. Inspired by the superpixel extraction methods [1, 40, 14, 29], we construct a soft feature element-region association matrix to make the clustering process differentiable.
@@
$\mathbf{D}(i,j)=\frac{{\exp}^{-\lVert e^{\prime\prime}_{i}-r^{\prime\prime}_{j}\rVert^{2}}}{\sum^{m}_{j=1}{\exp}^{-\lVert e^{\prime\prime}_{i}-r^{\prime\prime}_{j}\rVert^{2}}}$ (1)
@@
Note the feature regions are updated iteratively. Denote the current iteration as $t$. Once the region centers $\mathbf{R}^{\prime\prime}_{t}$ are updated, we can re-calculate the matrix $\mathbf{D}_{t+1}$ and then update $\mathbf{R}^{\prime\prime}_{t+1}$. This process is repeated until the maximum iteration $T$ is reached. The overview of this process is shown in Fig. 2 (right-top). The algorithm of this process is shown in Alg. 1 (Line 1 - Line 12).
3.1.2 Weighted Feature Aggregation for Node Representation
@@
Algorithm 1 Overall process of our method.
Input: Feature maps from the encoder $\{f^{1},...,f^{k}\}$, iteration number $T$
Output: Enhanced $\{f^{1},...,f^{k}\}$
1:  // In parallel
2:  for $f\in\{f^{1},...,f^{k}\}$ do
3:     Transforming $f$ to $f^{\prime\prime}$
4:     // Differentiable Feature Partition
5:     Initial region centers $\mathbf{R}^{\prime\prime}_{0}$
6:     for $t=1\;{\rm to}\;T$ do
7:        Calculating matrix $\mathbf{D}_{t}$ using Eq.(1),
8:        Updating region centers $\mathbf{R}^{\prime\prime}_{t}$ using Eq.(4),
9:     end for
10:     // Graph Nodes Formulation
11:     Obtaining the representations of graph nodes $\mathbf{R}$ using Eq.(5),
12:  end for
13:  // Hierarchical Graph Modeling
14:  Obtaining adjacency matrix $\mathbf{A}$,
15:  Obtaining graph nodes representations matrix $\mathbf{R}_{\rm G}$,
16:  // Structure-agnostic Graph Reasoning
17:  Updating $\mathbf{R}_{\rm G}$ by graph reasoning in Eq.(8),
18:  // Graph Features Inverse Mapping
19:  Mapping $\mathbf{R}_{\rm G}$ to $\{f^{1},...,f^{k}\}$ using Eq.(10),
20:  Fusing the mapped features with original features using Eq.(9) and Eq.(10).
3.1.3 Hierarchical Graph Modeling
@@
Since the regions corresponding to nodes are irregular, modeling the relationship is more challenging than the conventional grid-based methods. As such, we describe a 3D sliding-window based strategy to seek the correlations. It is noteworthy that in this step, we require an assignment of feature elements to regions. Therefore, we reuse Eq. (2) to divide each feature map into non-overlapped regions and assign each region a unique index from $\{1,...,M\}$. Then we can create the region index maps $\{J^{1},...,J^{k}\}$ corresponding to $\{f^{1},...,f^{k}\}$. We resize each index map to the same spatial size and concatenate them as a matrix $\mathbf{J}\in\mathbb{R}^{h\times w\times k}$. Then we perform the 3D sliding window with a size of $3\times 3\times 3$ in order. The criterion is that $\mathbf{A}(\mathbf{J}(p_{0}),\mathbf{J}(p_{0}+p_{n}))=1$, if $\mathbf{J}(p_{0})\neq\mathbf{J}(p_{0}+p_{n})$, where $p_{0}$ is the position of the window center and $p_{n}$ is the offset within window. To accelerate this process, we formulate it into CUDA code, see Supplementary for details.
Figure 3: Illustration of adjacent matrix $\mathbf{A}$. We build a hierarchical graph that establishes adjacency relationships within the same layer (intra-layer, orange arrows in Fig. 1 (b)) and also across different layers (inter-layer, purple arrows in Fig. 1 (b)), resulting in a single adjacent matrix $\mathbf{A}$. For simplicity, we use an example with two layers (highlighted in blue with one node and green with four nodes, see Fig 3 (a)). First, we resample them to the same resolution (see Fig 3 (b)). Then we stack them along channel dimension and apply the 3D sliding window strategy to obtain an adjacency matrix (see Fig 3 (c,d)).
@@
$\mathbf{R}_{\rm G}$ $=[\mathbf{R}^{1}_{\rm G};...;\mathbf{R}^{k}_{\rm G}]=[\mathbf{R}^{1}\mathbf{W}^{1};...;\mathbf{R}^{k}\mathbf{W}^{k}]$ (6)
@@
3.2 Structure-agnostic Graph Reasoning
In the design of our method, the feature regions can be dynamically updated during training, which subsequently results in different graph structures. To process the dynamic graphs, we adapt structure-agnostic graph reasoning strategies [9, 6, 38] to our method. Generally, this strategy consists of two steps: aggregation and update. The aggregation step gathers the information from surrounding nodes and the update step transforms the node representations based on the aggregated information with a learnable weight matrix.
@@ 3.2 Structure-agnostic Graph Reasoning
$\hat{\mathbf{R}}_{\rm{G}}=\left(\frac{1}{\mathbf{N_{A}}}\circ(\mathbf{A}\mathbf{R}_{\rm G})\right)\mathbf{W}_{\rm G}$ (7)
@@
To mitigate the over-smoothing effect, we employ a regularization on $\hat{\mathbf{R}}$ following [18, 10], which is defined as
@@
$\mathbf{F}^{i}_{\rm G}=\mathbf{D}^{i}\mathbf{R}^{i},\quad\mathbf{R}^{i}=\mathbf{R}^{i}_{\rm G}{\mathbf{W}^{i}_{\rm inv}}$ (9)
@@ 3.3 Objective Functions
4 Experiments
4.1 Experimental Settings
@@
Detection Decoder. To demonstrate the efficacy of our method, we only employ a plain feature pyramid detector [39]. The overview of the decoder is shown in Fig. 4 and the details are shown in Supplementary. The detection performance can potentially be further improved if a more advanced decoder is employed.
Figure 4: Architecture of the detection decoder.
[…]
Table 1: Comparison of different methods. The best result is highlighted in bold while the second-best result is marked by underlining.
@@
Figure 5: Qualitative visualization on CASIA, IMD20, NIST16, Coverage, and Columbia datasets (from top to bottom). More examples are in Supplementary
4.2 Comparing with State-of-the-arts
@@
Following previous methods, the comparison is conducted on CASIA [4], Coverage [36], and NIST16 [7] datasets since they provide training sets. Table 1 shows the results of our method compared to other state-of-the-art methods and graph-based methods under the fine-tuned scenario. We would like to clarify that EVP and NCL were trained without using synthetic dataset as others, while TransF. was tested on the Coverage dataset without fine-tuning. These three methods are marked by ()(*)( ∗ ). Moreover, the methods of PSCC-Net and HiFi-Net with ()($\dagger$)( † ) denote the results are reproduced by rigorously following their released codes. We have included the details of our reproduction in the Supplementary.
@@
4.3 Analysis
Ablation Study. This part investigates the effect of different settings: 1) Baseline indicates solely using Intern-tiny network. 2) Feature region $\rightarrow$ Grid indicates replacing the feature partition module with the original regular grid. 3) Graph reasoning $\rightarrow$ Linear refers to replacing graph reasoning with several linear layers (equivalent to fully connected graph reasoning). 4) Graph reasoning $\rightarrow$ MHSA refers to replacing graph reasoning with several Mutil-Head Self Attention [5] layers. 5) Hierarchical $\rightarrow$ Layer-wise means replacing the hierarchical graph modeling with layer-wise independent graph modeling and performing graph reasoning. 6) DFP w/o x- and y- coords means that we do not additionally concatenate the x and y coordinates in the clustering process of the DFP module. Table 2 displays the results of ablation experiments. It can be observed that when any module is replaced, the performance drops significantly, e.g., F1 scores decreasing by $2.9\%$, $2.6\%$, $3.1\%$, $1.9\%$, $1.0\%$ in the finetune stage. Nevertheless, only replacing one component still has an improvement compared to the baseline, which confirms the effectiveness of each component.
Table 2: Ablation study of different settings on CASIA dataset.
@@ 4.3 Analysis
Performance without Fine-tune. This scenario represents that the methods are pre-trained on synthetic images and directly tested on other datasets, aiming to exhibit the generalizability of detection. Table 3 shows the performance of the different base networks before and after using our method HRGR. These methods are pre-trained on the synthetic dataset from [21] and directly tested on all datasets. For a fair comparison, we use the same configuration for each experimental group. It can be seen that the performance of base networks generally improves after using our method.
Table 3: Performance without fine-tune.
@@
More Adaptations. To be more comprehensive, we conduct two more experiments in Table 4. The first part adapts the strategy of PSCC-Net into our backbone (Intern-tiny). The second adds HRGR to HRNet-based PSCC-Net. Both experiments use the same training configuration as in our method, suggesting that HRGR can be applied to different backbones and other IMD methods.
Table 4: More comparison with PSCC-Net on CASIA dataset.
Method CASIA Coverage NIST16 avg
AUC F1 AUC F1 AUC F1 AUC F1
Intern w/ PSCC 89.4 60.0 92.2 76.8 98.7 77.1 93.4 71.3
Intern w/ HRGR 93.4 69.9 94.8 82.1 99.3 84.6 95.8 78.9
PSCC w/o HRGR 89.4 59.6 93.1 70.0 95.8 60.7 92.7 63.4
PSCC w/ HRGR 89.9 61.3 92.0 70.4 96.6 64.2 92.8 65.3
Visualization of Feature Regions. We use the third layer of our model for demonstration. As shown in Fig. 6, our method generates non-overlapping and content-coherence feature regions, effectively delineating the boundary between manipulated and authentic regions, as highlighted in “GT matched feature regions”. In contrast, the feature regions within the authentic area (highlighted in gray) mostly appear as regular rectangles, indicating that our model has not overfitted to semantic features but instead concentrates on extracting manipulation traces.
Figure 6: Visualizations of feature partition. See text for details.
Attention Visualizations using GradGAM. In order to validate the effectiveness of the proposed Hierarchical Region-aware Graph Reasoning (HRGR), we present GradCAM [28] visualizations using the intermediate feature layer of the decoder, with (w/) and without (w/o) HRGR. Fig. 7 shows several examples. It can be observed that HRGR enables the achievement of more comprehensive and coherent attention to manipulated regions.
Figure 7: Attention visualizations without (w/o) and with (w/) HRGR using GradCAM.
Robustness. To evaluate the robustness of our method, we follow the perturbation settings in [13], which degrades manipulated images obtained from CASIA under different preprocesses, including resizing images with different scales, Gaussian Blurring with kernel size, adding Gaussian Noise with standard deviation “Sigma”, and performing JPEG Compression with quality factor. The evaluation results in comparison to PSCC-Net and HiFi-Net are shown in Fig. 8, which reveals that our method can resist certain perturbations. More analysis on other datasets can be found in Supplementary.
@@
Limitations. Since our method is performed on the encoders, its effectiveness depends on the capacities of these encoders. Our experiments have demonstrated that when using mainstream base networks, our method can further enhance the detection performance. Nevertheless, its effectiveness may be limited when applied to base networks with restricted capacities.
5 Conclusion

### v2

@@ Abstract
Image manipulation detection is to identify the authenticity of each pixel in images. One typical approach to uncover manipulation traces is to model image correlations. The previous methods commonly adopt the grids, which are fixed-size squares, as graph nodes to model correlations. However, these grids, being independent of image content, struggle to retain local content coherence, resulting in imprecise detection. To address this issue, we describe a new method named Hierarchical Region-aware Graph Reasoning (HRGR) to enhance image manipulation detection. Unlike existing grid-based methods, we model image correlations based on content-coherence feature regions with irregular shapes, generated by a novel Differentiable Feature Partition strategy. Then we construct a Hierarchical Region-aware Graph based on these regions within and across different feature layers. Subsequently, we describe a structural-agnostic graph reasoning strategy tailored for our graph to enhance the representation of nodes. Our method is fully differentiable and can seamlessly integrate into mainstream networks in an end-to-end manner, without requiring additional supervision. Extensive experiments demonstrate the effectiveness of our method in image manipulation detection, exhibiting its great potential as a plug-and-play component for existing architectures. Codes and models are available at https://github.com/OUC-VAS/HRGR-IMD.
Index Terms:
image manipualtion detection, feature clustering, graph reasoning
I Introduction
@@
Figure 1: The difference between (a) previous methods and (b) ours. Instead of using regular grids in previous methods, we model image correlations across different scales (orange and blue arrows) using connected local content-coherent feature regions.
To combat this, countermeasures have been proposed and gained increasing attention in recent years. Deep learning-based methods have emerged as the dominant method for detecting manipulations e.g., [2, 3, 4, 5, 6, 7],. Leveraging the impressive learning capabilities of deep learning models. Based on deep neural networks, feature enhancement strategies have been explored to further emphasize manipulation traces. Particularly, exploiting image correlations is a recent feature enhancement solution that has proven effective in enhancing detection ability [2, 3]. The intuition is that manipulated regions can be better perceived when compared to pristine regions. To capture the correlations, these methods usually employ the mechanism of either attention mechanism [3, 4] or graphs [8, 2]. For example, PSCC-Net [3] proposed a spatial-channel correlation module to capture the correlations using attention. ObjectFormer [4] learned the correlation between intermediate features in objects and patches through cross-attention. HGCN [8] and ERMPC [2] sought the relationship between pristine and manipulated regions by building a graph structure.
@@
II Method
II-A Hierarchical Region-aware Graph
@@
II-A1 Differentiable Feature Partition
@@
Intuitively, the conventional clustering algorithms (e.g., K-means) can be applied to feature layers by measuring the similarity among feature elements. However, this algorithm suffers from two difficulties. One is that local connectivity can not be guaranteed and the other is the clustering process is not differentiable. Inspired by the superpixel extraction methods [9, 10] and etc., we construct a soft feature element-region association matrix to make the clustering process differentiable.
@@
$\mathbf{D}(i,j)=\frac{{\exp}^{-\lVert e^{\prime\prime}_{i}-r^{\prime\prime}_{j}\rVert^{2}}}{\sum^{m}_{j=1}{\exp}^{-\lVert e^{\prime\prime}_{i}-r^{\prime\prime}_{j}\rVert^{2}}}.$ (1)
@@
Note the feature regions are updated iteratively. Denote the current iteration as $t$. Once the region centers $\mathbf{R}^{\prime\prime}_{t}$ are updated, we can re-calculate the matrix $\mathbf{D}_{t+1}$ and then update $\mathbf{R}^{\prime\prime}_{t+1}$. This process is repeated until the maximum iteration $T$ is reached. The overview of this process is shown in Fig. 2 (right-top).
II-A2 Weighted Feature Aggregation for Node Representation
@@
II-A3 Hierarchical Graph Modeling
@@
Since the regions corresponding to nodes are irregular, modeling the relationship is more challenging than the conventional grid-based methods. As such, we describe a 3D sliding-window based strategy to seek the correlations. It is noteworthy that in this step, we require an assignment of feature elements to regions. Therefore, we reuse Eq. (2) to divide each feature map into non-overlapped regions and assign each region a unique index from $\{1,...,M\}$. Then we can create the region index maps $\{J^{1},...,J^{k}\}$ corresponding to $\{f^{1},...,f^{k}\}$. We resize each index map to the same spatial size and concatenate them as a matrix $\mathbf{J}\in\mathbb{R}^{h\times w\times k}$. Then we perform the 3D sliding window with a size of $3\times 3\times 3$ in order. The criterion is that $\mathbf{A}(\mathbf{J}(p_{0}),\mathbf{J}(p_{0}+p_{n}))=1$, if $\mathbf{J}(p_{0})\neq\mathbf{J}(p_{0}+p_{n})$, where $p_{0}$ is the position of the window center and $p_{n}$ is the offset within window. To accelerate this process, we formulate it into CUDA code.
Figure 3: Illustration of adjacent matrix $\mathbf{A}$. We build a hierarchical graph that establishes adjacency relationships within the same layer (intra-layer, orange arrows in Fig. 1 (b)) and also across different layers (inter-layer, blue arrows in Fig. 1 (b)), resulting in a single adjacent matrix $\mathbf{A}$. For simplicity, we use an example with two layers (highlighted in blue with one node and green with four nodes, see Fig 3 (a)). First, we resample them to the same resolution (see Fig 3 (b)). Then we stack them along channel dimension and apply the 3D sliding window strategy to obtain an adjacency matrix (see Fig 3 (c,d)).
@@
$\mathbf{R}_{\rm G}$ $=[\mathbf{R}^{1}_{\rm G};...;\mathbf{R}^{k}_{\rm G}]=[\mathbf{R}^{1}\mathbf{W}^{1};...;\mathbf{R}^{k}\mathbf{W}^{k}],$ (6)
@@
II-B Structure-agnostic Graph Reasoning
In the design of our method, the feature regions can be dynamically updated during training, which subsequently results in different graph structures. To process the dynamic graphs, we adapt structure-agnostic graph reasoning strategies [11] and etc. to our method. Generally, this strategy consists of two steps: aggregation and update. The aggregation step gathers the information from surrounding nodes and the update step transforms the node representations based on the aggregated information with a learnable weight matrix.
@@ II-B Structure-agnostic Graph Reasoning
$\hat{\mathbf{R}}_{\rm{G}}=\left(\frac{1}{\mathbf{N_{A}}}\circ(\mathbf{A}\mathbf{R}_{\rm G})\right)\mathbf{W}_{\rm G},$ (7)
@@
To mitigate the over-smoothing effect, we employ a regularization on $\hat{\mathbf{R}}$ following [12], which is defined as
@@
$\mathbf{F}^{i}_{\rm G}=\mathbf{D}^{i}\mathbf{R}^{i},\quad\mathbf{R}^{i}=\mathbf{R}^{i}_{\rm G}{\mathbf{W}^{i}_{\rm inv}},$ (9)
@@ II-C Objective Functions
III Experiments
III-A Experimental Settings
@@
Implementation Details. Our method is implemented by PyTorch 11.3 with a Nvidia 3090 GPU. In the training phase, we utilize the AdamW optimizer with a learning rate of $6\times 10^{-5}$ and a weight decay of $0.05$. We employ a linear warm-up learning strategy, where the learning rate is linearly increased during warm-up and then linearly decreased based on the iteration until it reaches 00. The initial learning rate for warm-up is set to $1\times 10^{-6}$, and warm-up is completed within $1.5k$ iterations. The training image size is set to $512\times 512$ and the training iterations are set to $160k$. For feature integration, each layer’s $\mu_{i}$ is learned separately with initialization of $1$. In the objective function, we set the hyperparameters as follows: $\alpha=0.5,\gamma=2$. We employ a plain feature pyramid detector [20] as the decoder. The detection performance can be further improved if a more advanced decoder is employed.
TABLE I: Comparison of different methods. The best result is highlighted in bold while the second-best result is marked by underlining.
@@
Figure 4: Qualitative visualization on CASIA, IMD20, NIST16, Coverage, and Columbia datasets (from top to bottom).
III-B Comparing with State-of-the-arts
@@
Following previous methods, the comparison is conducted on CASIA, Coverage, and NIST16 datasets since they provide training sets. Table I shows the results of our method compared to other state-of-the-art methods and graph-based methods under the fine-tuned scenario. We would like to clarify that EVP and NCL were trained without using synthetic dataset as others, while TransF. was tested on the Coverage dataset without fine-tuning. These three methods are marked by ()(*)( ∗ ). Moreover, the methods marked with ()($\dagger$)( † ) are the ones that we have reproduced. Specifically, PSCC-Net and HiFi-Net were reproduced using the code provided by the respective authors, while the remaining methods were reproduced with IMDLBenCo [29].
@@
III-C Analysis
Ablation Study. This part investigates the effect of different settings: 1) Baseline indicates solely using Intern-tiny network. 2) Feature region $\rightarrow$ Grid indicates replacing the feature partition module with the original regular grid. 3) Graph reasoning $\rightarrow$ Linear refers to replacing graph reasoning with several linear layers (equivalent to fully connected graph reasoning). 4) Graph reasoning $\rightarrow$ MHSA refers to replacing graph reasoning with several Mutil-Head Self Attention layers. 5) Hierarchical $\rightarrow$ Layer-wise means replacing the hierarchical graph modeling with layer-wise independent graph modeling and performing graph reasoning. 6) DFP w/o x- and y- coords means that we do not additionally concatenate the x and y coordinates in the clustering process of the DFP module. Table II displays the results of ablation experiments. It can be observed that when any module is replaced, the performance drops significantly, e.g., F1 scores decreasing by $2.9\%$, $2.6\%$, $3.1\%$, $1.9\%$, $1.0\%$ in the finetune stage. Nevertheless, only replacing one component still has an improvement compared to the baseline, which confirms the effectiveness of each component.
TABLE II: Ablation study of different settings on CASIA dataset.
@@ III-C Analysis
Performance without Fine-tune. This scenario represents that the methods are pre-trained on synthetic images and directly tested on other datasets, aiming to exhibit the generalizability of detection. Table III shows the performance of the different base networks before and after using our method HRGR. These methods are pre-trained on the synthetic dataset from [3] and directly tested on all datasets. For a fair comparison, we use the same configuration for each experimental group. It can be seen that the performance of base networks generally improves after using our method.
TABLE III: Performance without fine-tune.
@@
Robustness. To evaluate the robustness of our method, we follow the perturbation settings in [3], which degrades manipulated images obtained from CASIA under different preprocesses, including resizing images with different scales, Gaussian Blurring with kernel size, adding Gaussian Noise with standard deviation “Sigma”, and performing JPEG Compression with quality factor. The evaluation results in comparison to PSCC-Net and HiFi-Net are shown in Fig. 5, which reveals that our method can resist certain perturbations.
@@
IV Conclusion